# Characterization of the directed landscape from the KPZ fixed point

### v2

@@ 5 Variants of the characterization theorem
In this section we prove Theorem 1.3, Corollary 1.5, and a technical variant of Theorem 1.3 that will be more appropriate for studying general AEPs. Throughout this section, we will appeal to the following simple lemma.
@@
and so since the right-hand side converges to 00 in probability, so does the left-hand side, yielding the conclusion. ∎
@@
Now, for $(x,s;y,t)\in\mathbb{Q}^{4}_{\uparrow}$ define $\mathcal{M}(x,s;y,t)=\mathcal{K}_{s,t}\delta_{x}(y)$. Then $\mathcal{M}$ satisfies conditions $1-3$ of Theorem 1.1. Condition $2$ follows by the independent increment property of $\mathcal{K}$, condition $3$ follows from (30), and for the triangle inequality (condition $1$), observe that for any rationals $q,q^{\prime}\in\mathbb{Q}$ with $q^{\prime}\leq\mathcal{K}_{s,r}\delta_{x}(y)\leq q$ we have that
@@
  • There exists a finite set $\{q_{1}<\dots<q_{k}\}\subset\mathbb{Q}$ such that $h$ is linear on each of the pieces $[q_{i},q_{i+1}]$, and $h(q_{i})\in\mathbb{Q}$ for all $i$.
  • $h^{\prime}=0$ off of the compact interval $[q_{1},q_{k}]$.
@@
In the remaining sections, we apply our general framework to prove convergence of specific models to the directed landscape. All convergence results are new, except for convergence of colored TASEP, colored ASEP and the KPZ equation. In those cases, our method offers a shorter route to proving convergence that bypasses a technical analysis of the underlying line ensembles.
@@ 6 Web distances
We next define these two models precisely. Let $\mathbb{Z}^{2}_{e}=\{(i,n)\in\mathbb{Z}^{2}:i+n\text{ is even}\}$ be the even lattice, and from every point $v=(i,n)$ assign two outgoing directed edges to $(i-1,n+1)$ and $(i+1,n+1)$. Let $\{\zeta(v),v\in\mathbb{Z}^{2}_{e}\}$ be a field of i.i.d. Rademacher random variables. The random walk web $Y$ is a family of coalescing random walks $Y_{(i,n)}:\llbracket n,\infty\rrbracket\to\mathbb{Z}$ started at each point $(i,n)\in\mathbb{Z}^{2}_{e}$, given by setting $Y_{(i,n)}(n)=i$ and
@@
    For all $i\geq 2$, given $B_{q_{1}},\dots,B_{q_{i-1}}$, let $T$ be the first time in $[s_{i},\infty)$ when $W_{q_{i}}(t)=B_{q_{j}}(t)$ for some $j<i$, and let $J$ be the minimal index such that $W_{q_{i}}(T)=B_{q_{J}}(T)$. Then set $B_{q_{i}}(t)=W_{q_{i}}(t)$ for $t\leq T$ and $B_{q_{i}}(t)=B_{q_{J}}(t)$ for $t\geq T$.
@@
Following [FINR04], we put a topology on the space $\Pi$ of all continuous functions $f:[t,\infty)\to\mathbb{R}$, where $t\in\mathbb{R}$. First, for such a function $f$, let $\hat{f}:\mathbb{R}\to\mathbb{R}$ be the function equal to $f$ on $[t,\infty)$ and equal to $f(t)$ on $(-\infty,t)$. We say that a sequence of continuous functions $f_{n}:[t_{n},\infty)\to\mathbb{R}$ converges to a limit $f:[t,\infty)\to\mathbb{R}$ if $t_{n}\to t$ and $\hat{f}_{n}\to\hat{f}$ uniformly on compact sets. This defines a Polish space. The Brownian web is the closure $\mathcal{B}$ of the rational Brownian web $\{B_{q}:q\in\mathbb{Q}^{2}\}$ in this space. Given the Brownian web, we can now define the Brownian web distance.
@@
such that $s<t_{1}<\cdots<t_{k}<t$, and there is a continuous path $\pi:[s,t]\to\mathbb{R}$ such that $\pi(t_{i})=x_{i}$ for each $i\in\llbracket 0,k+1,\rrbracket$, and for each $i\in\llbracket 0,k\rrbracket$ there exists $B_{i}\in\mathcal{B}$ with $\pi|_{[t_{i},t_{i+1}]}=B_{i}|_{[t_{i},t_{i+1}]}$. The infimum is equal to +$\infty$+ ∞ if there is no such $k\in\mathbb{N}$.
@@
~$\widetilde{\mathcal{M}}_{n}^{\eta}(0,0;y,t)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}(0,0;y,t)$
@@
The notations we use here are slight different from those in [VV23], but they are easily checked to be equivalent. Moreover, Theorem 1.9 in [VV23] (which concerns ~$\widetilde{\mathcal{M}}_{n}^{\eta}$) is stated only for $F=\{1\}$. However, the proof goes through for any finite set $F$ because of the coupling with the Seppäläinen-Johannson model/reflected simple random walks proven in that paper (see Proposition 3.7 therein). Vető and Virág chose to state their theorem only when $F=\{1\}$ only because they did not have access to the upgraded uniform convergence theorem (see the discussion after Theorem 1.9 in that paper).
@@
Let (~$(\widetilde{\mathcal{M}}_{n},n\in\mathbb{N})$ equal $(\mathcal{M}_{n}^{\eta},n\in\mathbb{N})$ for some fixed $\eta$, or $(\tilde{\mathcal{K}}_{n},n\in\mathbb{N})$. Then for any finite set $F\subset\mathbb{R}^{4}_{\uparrow}$, as $n\to\infty$, ~$\widetilde{\mathcal{M}}_{n}|_{F}\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}|_{F}.$
Given our Theorem 1.1, and Theorem 6.3, the only non-trivial task is to check that the triangle inequality passes to the limit. This is not obvious, since while $\mathcal{M}_{n}^{\eta},\mathcal{K}_{n}$ satisfy the triangle inequality, ~$\widetilde{\mathcal{M}}_{n}^{\eta},\widetilde{\mathcal{K}}_{n}$ do not.
@@
Fix $o=(x,s),p=(y,r),q=(z,t)\in\mathbb{R}^{2}$ with $s<r<t$. Let ~$\widetilde{\mathcal{M}}_{n}$ be as in Theorem 6.5. For any $\delta>0$, we have
@@
Let $\mathcal{M}_{n}$ equal $\mathcal{K}_{n}$ or $\mathcal{M}_{n}^{\eta}$, depending on whether ~$\widetilde{\mathcal{M}}_{n}$ equals $\widetilde{\mathcal{K}}_{n}$ or ~$\widetilde{\mathcal{M}}_{n}^{\eta}$. Then we have the metric composition law
@@
Let $g_{n}$ equal $g_{n}^{\eta}$ or $f_{n}$, to be narrow wedge used in the definition of ~$\widetilde{\mathcal{M}}_{n}$. Then if we let $y_{n}$ be the argmax of the map
@@
we have that ~$\widetilde{\mathcal{M}}_{n}(p,q)=g_{n}(y_{n}-y)+\mathcal{M}_{n}(y_{n},r;q)\leq\mathcal{M}_{n}(y_{n},r;q),$ and so by (37) we have that
~$\widetilde{\mathcal{M}}_{n}(o,p)+\widetilde{\mathcal{M}}_{n}(p,q)\leq\widetilde{\mathcal{M}}_{n}(o,p)+\mathcal{M}_{n}(y_{n},r;q)\leq\widetilde{\mathcal{M}}_{n}(o,q)+|\widetilde{\mathcal{M}}_{n}(o;p)-\widetilde{\mathcal{M}}_{n}(o;y_{n},r)|.$
Now, ~$\widetilde{\mathcal{M}}_{n}(o;\cdot,r)$ converges in law to a random continuous function uniformly on compact sets. This uses Theorem 6.3 and translation invariance of ~$\widetilde{\mathcal{M}}_{n}$, which follows from translation invariance of $D^{\operatorname{Br}}$ and $D^{\operatorname{RW}}$. Therefore to prove (36), it suffices to show that $y_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}y$ as $n\to\infty$.
@@
Let $Q\subset\mathbb{R}$ be a countable dense set such that $F\subset Q^{4}_{\uparrow}:=Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$. The sequence ~$\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$ is tight, since its one-point laws ~$\widetilde{\mathcal{M}}_{n}(u)$ are all tight by Theorem 6.3. Let ~$\widetilde{\mathcal{M}}:Q^{4}_{\uparrow}\to\mathbb{R}$ be any subsequential limit in law of ~$\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$. Then $\widetilde{\mathcal{M}}$ has independent increments, since the same was true of ~$\widetilde{\mathcal{M}}_{n}$. It also satisfies the triangle inequality by Lemma 6.6, and the marginals of ~$\widetilde{\mathcal{M}}(o;\cdot)$ match those of the directed landscape for any fixed $o\in Q^{2}$ by Theorem 6.3 and translation invariance of $\mathcal{M}_{n}$. Therefore by Theorem 1.1, ~$\widetilde{\mathcal{M}}=\mathcal{L}|_{Q^{4}_{\uparrow}}$ for some directed landscape $\mathcal{L}$, yielding the result. ∎
7 Polymer models
We next consider two polymer models: the KPZ equation and the O’Connell-Yor polymer. For the KPZ equation, convergence to the directed landscape was shown in [Wu23], and so our result here is simply an alternate proof. For the O’Connell-Yor polymer, the techniques used by [Wu23] to establish convergence of the KPZ equation should also be applicable, but the required technical work has not been carried out.
7.1 The KPZ equation
Introduced by Kardar, Parisi, and Zhang [KPZ86], the KPZ equation is the following stochastic PDE:
$\partial_{t}h=\frac{\lambda}{2}(\partial_{x}h)^{2}+\nu\partial^{2}_{x}h+\xi,$
where $\xi$ is a space-time Gaussian white noise, and $\lambda,\nu$ are parameters. Since the seminal work [BG97], a solution for the KPZ equation in terms of a Cole-Hopf transform has been widely used, and a more precise and rigorous solution theory was developed in [Hai13].
Over the past 15 years, exact formulas and long-time scaling limits have been proven for the KPZ equation, see e.g., [ACQ11, Dot13, SS10, IS11, BCFV15]. More recently, KPZ fixed point convergence was independently established in [QS23] and [Vir20], and directed landscape convergence was shown in [Wu23]. Here we give a short proof of directed landscape convergence, using the results of [QS23, Vir20] as the key inputs. Compared to [Wu23], our proof does not go through the KPZ line ensemble.
We start by stating KPZ fixed point convergence, following [QS23]. For $\delta>0$, consider the following KPZδ equation:
$\partial_{t}h=\frac{1}{4}(\partial_{x}h)^{2}+\frac{\delta}{4}\partial^{2}_{x}h+\delta^{1/2}\xi.$ (38)
We shall consider its Cole-Hopf solution $h=4\delta\log Z$, where $Z$ is the solution to the (rescaled) stochastic heat equation (SHE) with multiplicative noise:
$\partial_{t}Z=\delta\partial^{2}_{x}Z+\frac{1}{4}\delta^{-1/2}\xi Z.$ (39)
The solution for the SHE can be thought of as the partition function for a Brownian polymer weighted by the white noise $\xi$ [AKQ14]. More precisely, given the white noise $\xi$ there is a random field $Z^{\delta}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$ such that $Z^{\delta}(x,s;\cdot,\cdot)$ solves the SHE started from the Dirac $\delta$-measure at location $x$ and time $s$, where formally we have
[…]
Here the expectation is taken over a Brownian motion with $B(x)=s$ and variance $2\delta$. To make sense of the previous display rigorously and connect this with prelimiting models (i.e. the O’Connell-Yor polymer) one can use a chaos expansion. The SHE solution $Z$ for more general measure-valued initial conditions can be built from the field $Z^{\delta}$ using the fact that the SHE is linear in the initial condition. Applying the Cole-Hopf transform then solves the KPZ equation; here Dirac $\delta$-initial conditions for the SHE become narrow wedge initial conditions for the KPZ equation.
Theorem 7.1 ([QS23, Theorem 2.2(3)]).
Consider $f\in\operatorname{UC}$ which is either a sum of finitely many narrow wedge initial conditions, or is the maximum of such a function with a continuous function satisfying that $\sup_{x}f(x)/(1+|x|)<\infty$. Take the Cole-Hopf solution $h$ to (38) with initial condition $h(0,\cdot)=f$. As $\delta\to 0$ we have $h(t,\cdot)+\frac{t}{12}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$, in the uniform-on-compact topology, jointly for finitely many $t>0$. In particular, this holds for $h(t,\cdot)=4\delta\log Z^{\delta}(x,s;\cdot,s+t)$ (the narrow wedge initial condition at location $x$, time $s$).
Using Theorem 1.1 and Theorem 7.1, we can recover the convergence of the KPZ equation to the directed landscape from [Wu23].
Theorem 7.2.
The function $(x,s;y,t)\mapsto\mathcal{H}^{\delta}(x,s;y,t)+\frac{t-s}{12}$ converges in law to $\mathcal{L}$ as $\delta\to 0$, in the sense of finite dimensional distributions.
We remark that this finite dimensional convergence can be upgraded to (1) uniform-on-compact convergence, or (2) the convergence of multiple KPZ equations with initial conditions as in Theorem 7.1 coupled via the same space-time white noise. For these stronger convergence statements, see [Wu23, Theorem 1.6, Corollary 1.9]. To derive these upgrades from Theorem 7.2, it suffices to prove the tightness of $\mathcal{H}^{\delta}$ as a continuous function on $\mathbb{R}^{4}_{\uparrow}$. This follows from modulus of continuity bounds of $\mathcal{H}^{\delta}$ from e.g., [CGH21, DG23], and has been detailed in [Wu23, Proposition 6.2].
Proof of Theorem 7.2.
Fix a finite set $F\subset\mathbb{R}^{4}_{\uparrow}$, and let $Q\subset\mathbb{R}$ be a countable dense set such that $F\subset Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$. Let $\mathcal{M}:Q^{4}\cap\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$ be a joint subsequential limit in law as $\delta\to 0$ of $(x,s;y,t)\mapsto\mathcal{H}^{\delta}(x,s;y,t)+\frac{t-s}{12},(x,s;y,t)\in Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$.
We will apply Theorem 1.1 to $\mathcal{M}$. We check that the three conditions of that theorem are satisfied. Independence of increments is obvious, and the KPZ fixed point marginal condition follow from Theorem 7.1. We next prove the triangle inequality. Take any $o=(x,s)$, $p=(y,r)$, $q=(z,t)\in Q^{2}$ with $s<r<t$. The
$\exp\left((4\delta)^{-1}\mathcal{H}^{\delta}(o;q)\right)$ $=\int\exp\left((4\delta)^{-1}(\mathcal{H}^{\delta}(o;w,r)+\mathcal{H}^{\delta}(w,r;q))\right)\mathrm{d}w$
$\geq\int_{y-\delta}^{y+\delta}\exp\left((4\delta)^{-1}(\mathcal{H}^{\delta}(o;w,r)+\mathcal{H}^{\delta}(w,r;q))\right)\mathrm{d}w.$
Here the first line is the Chapman-Kolmogorov equation for the SHE (e.g. see [AKQ14, Theorem 3.1]); this should be viewed as a positive temperature version of the metric composition law. In the KPZ scaling limit as $\delta\to 0$, it becomes the usual metric composition law. In particular, the above computation implies that
$\mathcal{H}^{\delta}(o;q)\geq$ $\mathcal{H}^{\delta}(o;p)+\mathcal{H}^{\delta}(p;q)+4\delta\log(2\delta)$
$+\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(o;w,r)-\mathcal{H}^{\delta}(o;p)+\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q).$
By Theorem 7.1, we have
$\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(o;w,r)-\mathcal{H}^{\delta}(o;p)\to 0,\quad\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q)\to 0$
in probability as $\delta\to 0$. For the second convergence, we used that $\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q)$ has the same law as $\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(-q;-w,-r)-\mathcal{H}^{\delta}(-q;-p)$, due to the reflection symmetry of $\mathcal{H}^{\delta}$ (see e.g., [AJRAS22, Proposition 2.3]). Also note that $4\delta\log(2\delta)\to 0$ as $\delta\to 0$. Therefore we have $\mathcal{M}(o;q)\geq\mathcal{M}(o;p)+\mathcal{M}(p;q)$, i.e.  the triangle inequality holds. Then by Theorem 1.1, $\mathcal{M}$ has the same law as $\mathcal{L}$ restricted to $Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$, so the conclusion follows. ∎
7.2 O’Connell-Yor polymer
The O’Connell-Yor polymer [OY01] is an integrable semi-discrete directed polymer model defined through a family of independent Brownian motions. See e.g., [O’C12, BCF14, IS16, Nic21, Vir20, LNS23] and references therein for a selection of different perspectives on the model. In particular, Borodin, Corwin, and Ferrari [BCF14] proved convergence of the one-point distributions of the O’Connell-Yor polymer starting from a certain family of initial conditions corresponding to the Baik-Ben Arous-Peché statistics from random matrix theory. Virág [Vir20] built on these results to show that the O’Connell-Yor polymer converges to the KPZ fixed point from an arbitrary initial condition. We use Theorem 1.3 to upgrade Virág’s theorem to give directed landscape convergence.
To define the O’Connell-Yor polymer, consider a collection $B=\{B_{i}\}_{i\in\mathbb{Z}}$ of independent two-sided Brownian motions. In this section only, our Brownian motions have variance $1$ rather than variance $2$ to match the conventions in the literature. For $n<m\in\mathbb{Z}$ and $x<y\in\mathbb{R}$, define the polymer free energy at inverse temperature $\beta>0$ by
$H^{\beta}_{n,m}(x,y)=\beta^{-1}\log\int_{x=x_{n}<x_{n+1}<\dots<x_{m}=y}\mathrm{d}x_{n+1}\cdots\mathrm{d}x_{m-1}\exp\left(\sum_{i=n+1}^{m}\beta[B_{i}(x_{i})-B_{i}(x_{i-1})]\right).$
We also let $H^{\beta}_{n,m}(x,y)=-\infty$ when $n<m\in\mathbb{Z}$ and $x\geq y$ or when $n=m\in\mathbb{Z}$ and $x\neq y$, and set $H^{\beta}_{n,n}(x,x)=0$ for any $n\in\mathbb{Z}$ and $x\in\mathbb{R}$.
We will consider limits of the free energy as the distance between $(x,n),(y,m)\to\infty$ along a fixed diagonal in the plane. From now on we will assume $\beta=1$ as the corresponding results for other $\beta$ can be recovered by applying a shear transformation of the model and Brownian scaling. Following [BCF14], let $\Psi(\theta)=(\log\Gamma(\theta))^{\prime}$ be the digamma function and define functions $a=a(\theta),b=b(\theta),c=c(\theta)$ by
$a=-2/\Psi^{\prime\prime},\qquad b=a\Psi^{\prime},\qquad c=\theta b-a\Psi.$
Below we fix $\theta>0$, and for $s\leq t\in(an^{3})^{-1}\mathbb{Z}$ and $x,y\in\mathbb{R}$, set
$\mathcal{H}^{n}(x,s;y,t)=n^{-1}\left(H^{1}_{asn^{3},atn^{3}}(bsn^{3}+2xn^{2},btn^{3}+2yn^{2})-c(t-s)n^{3}+2(x-y)n^{2}\Psi^{\prime}\right).$
For any other $s\leq t\in\mathbb{R}$, we let
$\mathcal{H}^{n}(x,s;y,t)=\mathcal{H}^{n}(x,(an^{3})^{-1}\lfloor asn^{3}\rfloor;y,(an^{3})^{-1}\lfloor atn^{3}\rfloor),$
and define $\mathcal{H}^{n}_{s,t}=\mathcal{H}^{n}(\cdot,s;\cdot,t)$. As with the KPZ equation, the field $\mathcal{H}^{n}_{s,t},s\leq t\in\mathbb{R}$ satisfies a positive temperature version of the metric composition law:
$\mathcal{H}^{n}_{s,t}(x,y)=n^{-1}\log\int\exp\left(n\mathcal{H}^{n}_{s,r}(x,z)+n\mathcal{H}^{n}_{r,t}(z,y)\right)\mathrm{d}z,$ (40)
for any $s\leq r\leq t\in\mathbb{R}$ and $x,y\in\mathbb{R}$.
Theorem 7.3 ([Vir20], Corollary 12).
For any fixed $s<t$, $\mathcal{H}^{n}_{s,t}$ is tight in the uniform-on-compact topology as $n\to\infty$. Moreover, any subsequential limit $\mathcal{H}^{\infty}_{s,t}$ is continuous with KPZ fixed point marginals: for any $f,g\in\operatorname{UC}$, $f\diamond\mathcal{H}^{\infty}_{s,t}\diamond g\stackrel{{\scriptstyle d}}{{=}}f\diamond\mathcal{L}_{s,t}\diamond g$.
We can upgrade this KPZ fixed point convergence to the following directed landscape convergence.
Theorem 7.4.
We have $\mathcal{H}^{n}\to\mathcal{L}$ as $n\to\infty$, in the sense that for any fixed $s_{1}<t_{1},\ldots,s_{k}<t_{k}$, we have $\{\mathcal{H}^{n}_{s_{i},t_{i}}\}_{i=1}^{k}\stackrel{{\scriptstyle d}}{{\to}}\{\mathcal{L}_{s_{i},t_{i}}\}_{i=1}^{k}$ in the uniform in compact topology.
Proof.
Let $Q\subset\mathbb{R}$ be a countable dense set containing $s_{i},t_{i}$ for each $i\in\llbracket 1,k\rrbracket$. By Theorem 7.3, $\{\mathcal{H}^{n}_{s,t}\}_{s<t\in Q}$ is tight in the uniform-on-compact topology. To prove the theorem, it suffices to show that any limit point $\{\mathcal{H}^{\infty}_{s,t}\}_{s<t\in Q}$ has the same distribution as $\{\mathcal{L}_{s,t}\}_{s<t\in Q}$. For the convenience of notations, below we denote $\mathcal{H}^{\infty}_{s,s}(x,x)=0$ and $\mathcal{H}^{\infty}_{s,s}(x,y)=-\infty$ for any $s\in Q$ and $x\neq y$.
We will apply Theorem 1.3. For this, define a family of random operators $\{\mathcal{K}_{s,t}\}_{s\leq t\in Q}$ on $\operatorname{NW}$, as follows. For $s\leq t\in Q$, and $f\in\operatorname{NW}$, we let $\mathcal{K}_{s,t}f=f\diamond\mathcal{H}^{\infty}_{s,t}$.
The operators $\mathcal{K}_{s,t}$ have KPZ fixed point marginals by Theorem 7.3, the independent increment property is inherited from the O’Connell-Yor polymer, and shift commutativity are inherited from the corresponding properties of the O’Connell-Yor polymer. It then remains to check monotonicity, i.e., condition 3 of Theorem 1.3.
From (40) and Theorem 7.3, and using the same arguments as in the proof of Theorem 7.2, we have that for any $o=(x,s)$, $p=(y,r)$, $q=(z,t)\in\mathbb{R}\times Q$ with $s<r<t$,
$\mathcal{H}^{\infty}_{s,r}(x,y)+\mathcal{H}^{\infty}_{r,t}(y,z)\leq\mathcal{H}^{\infty}_{s,t}(x,z).$ (41)
Then for any $f,g\in\operatorname{NW}$ and $s\leq r<t\in Q$ with $g\leq\mathcal{K}_{s,r}f$, we have
$\mathcal{K}_{r,t}g=g\diamond\mathcal{H}^{\infty}_{r,t}\leq\mathcal{K}_{s,r}f\diamond\mathcal{H}^{\infty}_{r,t}=f\diamond\mathcal{H}^{\infty}_{s,r}\diamond\mathcal{H}^{\infty}_{r,t}\leq f\diamond\mathcal{H}^{\infty}_{s,t}=\mathcal{K}_{s,t}f,$
where the second inequality is due to the triangle inequality (41). Thereby we get monotonicity, as stated in condition 3 of Theorem 1.3.
Now by Theorem 1.3, we conclude that there exists a directed landscape $\mathcal{L}$, such that for any $s<t\in Q$ and $x\in Q$, we have
$\mathcal{H}^{\infty}_{s,t}(x,\cdot)=\mathcal{L}_{s,t}(x,\cdot).$
As both $\mathcal{H}^{\infty}_{s,t}$ and $\mathcal{L}_{s,t}$ are continuous, we necessarily have that $\mathcal{H}^{\infty}_{s,t}=\mathcal{L}_{s,t}$. The conclusion follows. ∎

### v3

@@ 5 Variants of the characterization theorem
In this section we prove Theorem 1.3, Corollary 1.5, and a technical variant of Theorem 1.3 that will be used later to study convergence of general exclusion processes. Throughout this section, we will appeal to the following simple lemma.
@@
and so since the right-hand side converges to 0 in probability, so does the left-hand side, yielding the conclusion. ∎
@@
Now, for $(x,s;y,t)\in\mathbb{Q}^{4}_{\uparrow}$ define $\mathcal{M}(x,s;y,t)=\mathcal{K}_{s,t}\delta_{x}(y)$. Then $\mathcal{M}$ satisfies conditions 131-3 of Theorem 1.1. Condition $2$ follows by the independent increment property of $\mathcal{K}$, condition $3$ follows from (31), and for the triangle inequality (condition $1$), observe that for any rationals $q,q^{\prime}\in\mathbb{Q}$ with $q^{\prime}\leq\mathcal{K}_{s,r}\delta_{x}(y)\leq q$ we have that
@@
  • There exists a finite set $\{q_{1}<\dots<q_{k}\}\subset\mathbb{Q}$ such that $h$ is linear on each of the pieces $q_{i},q_{i+1}]$, and $h(q_{i})\in\mathbb{Q}$ for all $i$.
  • $h^{\prime}=0$ off of the compact interval $q_{1},q_{k}]$.
@@
In the remaining sections, we apply our general framework to prove convergence of specific models to the directed landscape. All convergence results are new, except for convergence of colored TASEP and ASEP. In the case of colored TASEP, our method offers a shorter route to proving convergence that bypasses a technical analysis of the underlying line ensembles.
@@ 6 Web distances
We next define these two models precisely. Let $\mathbb{Z}^{2}_{e}=\{(i,n)\in\mathbb{Z}^{2}:i+n\text{ is even}\}$ be the even lattice, and from every point $v=(i,n)$ assign two outgoing directed edges to (i1,n+1)(i-1,n+1) and $(i+1,n+1)$. Let $\{\zeta(v),v\in\mathbb{Z}^{2}_{e}\}$ be a field of i.i.d. Rademacher random variables. The random walk web $Y$ is a family of coalescing random walks $Y_{(i,n)}:\llbracket n,\infty\rrbracket\to\mathbb{Z}$ started at each point $(i,n)\in\mathbb{Z}^{2}_{e}$, given by setting $Y_{(i,n)}(n)=i$ and
@@
    For all $i\geq 2$, given $B_{q_{1}},\dots,B_{q_{i-1}}$, let $T$ be the first time in $s_{i},\infty)$ when $W_{q_{i}}(t)=B_{q_{j}}(t)$ for some $j<i$, and let $J$ be the minimal index such that $W_{q_{i}}(T)=B_{q_{J}}(T)$. Then set $B_{q_{i}}(t)=W_{q_{i}}(t)$ for $t\leq T$ and $B_{q_{i}}(t)=B_{q_{J}}(t)$ for $t\geq T$.
@@
Following [FINR04], we put a topology on the space $\Pi$ of all continuous functions $f:[t,\infty)\to\mathbb{R}$, where $t\in\mathbb{R}$. First, for such a function $f$, let $\hat{f}$$\mathbb{R}\to\mathbb{R}$ be the function equal to $f$ on $t,\infty)$ and equal to $f(t)$ on $(-\infty,t)$. We say that a sequence of continuous functions $f_{n}:[t_{n},\infty)\to\mathbb{R}$ converges to a limit $f:[t,\infty)\to\mathbb{R}$ if $t_{n}\to t$ and $\hat{f}_{n}\to\hat{f}$ uniformly on compact sets. This defines a Polish space. The Brownian web is the closure $\mathcal{B}$ of the rational Brownian web $\{B_{q}:q\in\mathbb{Q}^{2}\}$ in this space. Given the Brownian web, we can now define the Brownian web distance.
@@
such that $s<t_{1}<\cdots<t_{k}<t$, and there is a continuous path $\pi:[s,t]\to\mathbb{R}$ such that $\pi(t_{i})=x_{i}$ for each $i\in\llbracket 0,k+1,\rrbracket$, and for each $i\in\llbracket 0,k\rrbracket$ there exists $B_{i}\in\mathcal{B}$ with $\pi|_{[t_{i},t_{i+1}]}=B_{i}|_{[t_{i},t_{i+1}]}$. The infimum is equal to $\infty$ if there is no such $k\in\mathbb{N}$.
@@
$\widetilde{\mathcal{M}}_{n}^{\eta}(0,0;y,t)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}(0,0;y,t)$
@@
The notations we use here are slight different from those in [VV23], but they are easily checked to be equivalent. Moreover, Theorem 1.9 in [VV23] (which concerns $\widetilde{\mathcal{M}}_{n}^{\eta}$) is stated only for $F=\{1\}$. However, the proof goes through for any finite set $F$ because of the coupling with the Seppäläinen-Johannson model/reflected simple random walks proven in that paper (see Proposition 3.7 therein). Vető and Virág chose to state their theorem only when $F=\{1\}$ only because they did not have access to the upgraded uniform convergence theorem (see the discussion after Theorem 1.9 in that paper).
@@
Let $(\widetilde{\mathcal{M}}_{n},n\in\mathbb{N})$ equal $(\mathcal{M}_{n}^{\eta},n\in\mathbb{N})$ for some fixed $\eta$, or $(\tilde{\mathcal{K}}_{n},n\in\mathbb{N})$. Then for any finite set $F\subset\mathbb{R}^{4}_{\uparrow}$, as $n\to\infty$, $\widetilde{\mathcal{M}}_{n}|_{F}\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}|_{F}.$
Given our Theorem 1.1, and Theorem 6.3, the only non-trivial task is to check that the triangle inequality passes to the limit. This is not obvious, since while $\mathcal{M}_{n}^{\eta},\mathcal{K}_{n}$ satisfy the triangle inequality, $\widetilde{\mathcal{M}}_{n}^{\eta},\widetilde{\mathcal{K}}_{n}$ do not.
@@
Fix $o=(x,s),p=(y,r),q=(z,t)\in\mathbb{R}^{2}$ with $s<r<t$. Let $\widetilde{\mathcal{M}}_{n}$ be as in Theorem 6.5. For any $\delta>0$, we have
@@
Let $\mathcal{M}_{n}$ equal $\mathcal{K}_{n}$ or $\mathcal{M}_{n}^{\eta}$, depending on whether $\widetilde{\mathcal{M}}_{n}$ equals $\widetilde{\mathcal{K}}_{n}$ or $\widetilde{\mathcal{M}}_{n}^{\eta}$. Then we have the metric composition law
@@
Let $g_{n}$ equal $g_{n}^{\eta}$ or $f_{n}$, to be narrow wedge used in the definition of $\widetilde{\mathcal{M}}_{n}$. Then if we let $y_{n}$ be the argmax of the map
@@
we have that $\widetilde{\mathcal{M}}_{n}(p,q)=g_{n}(y_{n}-y)+\mathcal{M}_{n}(y_{n},r;q)\leq\mathcal{M}_{n}(y_{n},r;q),$ and so by (38) we have that
$\widetilde{\mathcal{M}}_{n}(o,p)+\widetilde{\mathcal{M}}_{n}(p,q)\leq\widetilde{\mathcal{M}}_{n}(o,p)+\mathcal{M}_{n}(y_{n},r;q)\leq\widetilde{\mathcal{M}}_{n}(o,q)+|\widetilde{\mathcal{M}}_{n}(o;p)-\widetilde{\mathcal{M}}_{n}(o;y_{n},r)|.$
Now, $\widetilde{\mathcal{M}}_{n}(o;\cdot,r)$ converges in law to a random continuous function uniformly on compact sets. This uses Theorem 6.3 and translation invariance of $\widetilde{\mathcal{M}}_{n}$, which follows from translation invariance of $D^{\operatorname{Br}}$ and $D^{\operatorname{RW}}$. Therefore to prove (37), it suffices to show that $y_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}y$ as $n\to\infty$.
@@
Let $Q\subset\mathbb{R}$ be a countable dense set such that $F\subset Q^{4}_{\uparrow}:=Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$. The sequence $\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$ is tight, since its one-point laws $\widetilde{\mathcal{M}}_{n}(u)$ are all tight by Theorem 6.3. Let $\widetilde{\mathcal{M}}:Q^{4}_{\uparrow}\to\mathbb{R}$ be any subsequential limit in law of $\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$. Then $\widetilde{\mathcal{M}}$ has independent increments, since the same was true of $\widetilde{\mathcal{M}}_{n}$. It also satisfies the triangle inequality by Lemma 6.6, and the marginals of $\widetilde{\mathcal{M}}(o;\cdot)$ match those of the directed landscape for any fixed $o\in Q^{2}$ by Theorem 6.3 and translation invariance of $\mathcal{M}_{n}$. Therefore by Theorem 1.1, $\widetilde{\mathcal{M}}=\mathcal{L}|_{Q^{4}_{\uparrow}}$ for some directed landscape $\mathcal{L}$, yielding the result. ∎